# AlphaEarth Foundations: An embedding field model for accurate and efficient global mapping from sparse label data

### v1

@@ Abstract
Unprecedented volumes of Earth observation data are continually collected around the world, but high-quality labels remain scarce given the effort required to make physical measurements and observations. This has led to considerable investment in bespoke modeling efforts translating sparse labels into maps. Here we introduce AlphaEarth Foundations, an embedding field model yielding a highly general, geospatial representation that assimilates spatial, temporal, and measurement contexts across multiple sources, enabling accurate and efficient production of maps and monitoring systems from local to global scales. The embeddings generated by AlphaEarth Foundations are the only to consistently outperform all previous featurization approaches tested on a diverse set of mapping evaluations without re-training. We will release a dataset of global, annual, analysis-ready embedding field layers from 2017 through 2024.
@@ 2 From sparse labels to maps
evaluation shown, all available training data or the "max-trial" was used. Error bars indicate the $1\sigma$ best and worst case ratio or $\sim$90confidence interval by bootstrapping and k-folds when possible. Evaluations are sorted by increasing mean error ratio. The dashed line represents a ratio of 1 with higher values indicating AEF outperforming the next best solution. (B) AEF reconciles multiple sparse, non-uniformly sampled observation records into a continuous record, regardless of fluctuations in availability. AEF "embedding fields" are a result of static temporal summaries drawn over a conditional "valid period" that need not fully intersect the "support period", where the latter defines the temporal range of the input data. Multiple raster and scalar measurement sources are modeled as sources or targets by AEF. These may be any combination of temporally, geographically, and spatially sparse. In the example shown here, NLCD is not present, and GEDI is available in only a sparse fraction of the spatial context. (C) A view of our global embedding field for the year 2023, note apparent climatic gradients at large scales. (D) AEF produces highly resolved features at 10m2, shown here plotting arbitrary axes in Oaxaca, Mexico. (E) A stack of 64 rasterized AEF layers forms an embedding field, and each individual vector maps to a coordinate on the unit sphere $S^{63}$.
@@ 3 AlphaEarth Foundations
AlphaEarth Foundations (AEF) is the only task-agnostic learned EO featurization approach to outperform all existing featurization methods tested across a broad set of sparse data domains (Figure 1A), reducing error magnitudes by $\sim$23.9($\sim$1.4x error magnitude reduction) on average while maintaining a best-in-class 10-meter spatial resolution that requires 16x less information per-representation (64 bytes) compared to the next-most compact learned method. We achieve this leap in performance across challenging mapping applications through a number of innovations; namely, we employ an adaptive decoding scheme that considers time and sensor parameters as continuous variables in an implicit decoder with associated losses, a spatially dense information time-bottleneck, time-conditional summarization, and spatially-precise alignment with geotagged text. To the best of our knowledge, AEF is the first EO featurization approach to support continuous time (Figure 1B; see supplemental materials S16.2.1). Additionally we introduce a challenging evaluation suite composed of high-quality reference data that attempts to faithfully replicate realistic mapping scenarios. We make our annualized planet-scale feature maps or "embedding field" layers (Figure 1C-D) and evaluation suite available under an open license to encourage further exploration and use.
@@
AEF is designed to accept $N_{i}$ frames for $i\in M_{E}$ input (encoded) data sources with $C_{i}$ channels resampled to the same spatial resolution, and a millisecond epoch timestamp $t_{j},1\leq j\leq{\Sigma}N_{i}$ (Figure 2A). The range of the input timestamps we refer to as the “support period”. For the purposes of learning or at inference time, we support a pair of conditioning timestamps or “valid period” $t_{s},t_{e}$ where $t_{s}<t_{e}$ provides a temporal summary of the Earth’s surface and climatic activity over $[t_{s},t_{e})$, even when there is no $t_{j}$ for which $t_{s}\leq t_{j}<t_{e}$ (interpolation), or when $t_{e}\leq t_{j},\forall j\in\{1,2,...,{\Sigma}N_{i}\}$ or $t_{s}>t_{j},\forall j\in\{1,2,...,{\Sigma}N_{i}\}$ (extrapolation). These summaries or “embeddings” are 64 bytes in size, and each embedding contains information that reproduces the temporal trajectory of variables listed in Table S1 over the summary period (Figure 2B) using conditional metadata from each source (see supplemental materials S16.2.1). By explicitly separating the input intervals from those used for the temporal summary, we can apply AEF to time dependent problems requiring a precise date range without fine-tuning. Embeddings are further constrained to distribute uniformly in $S^{63}$ using a so-called “batch uniformity” objective (Figure 2C; see supplemental materials S16.2.4).
@@
Figure 2: AlphaEarth Foundations. (A) Block diagram of the overall network architecture used for video analysis. Preprocessing converts raw observation data via normalization using global statistics, and acquisition timestamps are converted to sinusoidal timecodes. Individual source encoders transform inputs to the same latent space before entering the bulk of the model. Outputs are summarized using conditional timecodes or "summary periods", unique to each decoded source and contrastive learning task. $\mu$ refers to the embedding outputs of the model. (B) Model outputs are treated as the mean direction of a von Mises-Fisher distribution, and decoding proceeds by sampling this distribution, and concatenating it with sensor geometry metadata and a timecode indicating the relative position in the valid period to decode. Decoding proceeds for all sources, with losses dependent on the characteristics of each source (see supplemental materials S15). (C) To prevent collapse and improve performance, embeddings are compared to equivalent batch-rotated embeddings using a dot product. The absolute value of this quantity is minimized as a necessary condition for an empirically uniform distribution in $S^{63}$. (D) Block diagram of the model bulk, consisting of simultaneous pathways at different resolutions to maintain efficiency and spatial precision. (E) Contrastive learning between the video teacher and student model, and text encoder. (F) Complete 360° view of 2023 annual embedding field covering Earth’s land surface including minor islands over approximately $\pm$ 82°.
@@
Our video summarization architecture must simultaneously maintain highly localized representations as well as model long distance relationships through time and space in a computationally efficient way; for this we’ve designed an encoder termed Space Time Precision or “STP” that consists of repeated blocks of three simultaneous operators interleaved with spatial pyramid "exchanges" (Figure 2D) inspired by Wang et al. (2020) but more efficiently utilizing learned resampling stages. Given a square input of $L$ pixels a side, each block consists of a $\frac{1}{16}L$space" operator following ViT-like spatial self-attention (Dosovitskiy et al., 2020), a $\frac{1}{8}L$ “time” operator utilizing time-axial self-attention, and a $\frac{1}{2}L$precision" operator utilizing 3x3 convolutions. Each sequence element in the “time” operator is conditioned on the associated input timestamp ($t_{j}$) after conversion to a sinusoidal timecode. STP blocks terminate with learned Laplacian pyramid rescaling, such that each operator can pass its state to each of the operators in the subsequent block. STP itself terminates with a final learned spatial resampling to the resolution of the "precision" operator. Thus for ${\sum}N_{i}$ inputs, STP produces ${\sum}N_{i}$ output feature maps at $\frac{1}{2}L$ pixels resolution.
@@ 4 Evaluation in realistic data-scarce scenarios
To establish the performance of AEF relative to other domain-specific learned and designed representations, we required an evaluation dataset that included archetypal examples of realistic mapping applications. These include thematic mapping, biophysical variable estimation, and change detection at annual and sub-annual cadences. We found most if not all datasets in existing geospatial benchmark suites provide annotations at an object- or image-level rather than pixel-level, rely on labels sampled from existing (machine-generated) datasets, require running the benchmark analysis using provided source imagery, have limited geographic coverage, and/or do not provide sufficient spatial precision or temporal information, e.g., (Lacoste et al., 2023; Bountos et al., 2023), limiting their value in assessing practical use.
To address the need for high-quality labeled datasets that can be used to simulate low-shot (i.e., tens to hundreds of samples) performance in data-scarce regimes requiring precise map outputs, we developed a set of 15 evaluations sourced from 11 openly available datasets representing land use/land cover mapping and change detection, crop type mapping at different hierarchies, tree genera and plantation classification, and estimating evapotranspiration and emissivity (Table 1). For each evaluation dataset, we selected a balanced number of training samples from each class or equally-spaced partition, with the number of samples determined based on the minimum class size (Table 1, Max Trial Size) and the remainder of the dataset reserved for testing. We then assess performance across a suite of trials designed to test very-low shot training sizes with $n$ samples per class ($n=1$, 101010, max) via transferal methods with minimal parameters: k-nearest neighbors, and linear layers fit to the features (see supplemental materials S18). The “max-trial” scenario is meant to represent realistic dataset sizes whereas the ten and one shot trials are meant to evaluate performance given extreme data sparsity.
@@
We used this set of evaluations to compare AEF with domain-specific baselines including three designed featurization approaches: CCDC (Zhu and Woodcock, 2014; Gorelick et al., 2023), MOSAIKS (Rolf et al., 2021), and composites (Qiu et al., 2023), and three learned featurization approaches: SatCLIP (Klemmer et al., 2025), Prithvi (Jakubik et al., 2023b), and Clay (Clay, 2024). We also include three controls: spatial coordinates (XY), coordinates and elevation (XYZ), and a ViT (Vision Transformer) pre-trained on ImageNet (Dosovitskiy et al., 2020; Deng et al., 2009). Where applicable, baselines were provided with identical inputs to AEF and baseline hyperparameters were tuned to maximize performance on our evaluation suite (see supplemental materials S19).
Our evaluations showed AEF consistently outperforms designed and learned featurization methods in all trial settings. AEF reduced error magnitudes overall by $\sim$23.9on average when compared to the next-best approach and method of transfer in the max-trial setting (Figure 1A). For ten-shot trials, AEF reduced error magnitudes by $\sim$10.4on average compared to the next-best approach, and for one-shot trials AEF reduced error magnitudes by $\sim$4.18%. We show quantitative and qualitative results for select evaluations in Figure 3 (and see supplemental materials S20 for full quantitative results). The next-best approach varies across evaluation dataset and method, indicating both non-uniform progress and that AEF unlocks progress in historically challenging mapping scenarios. We discuss these results and the effect of scaling training data in further detail below.
@@ 5 Thematic mapping
Thematic mapping or "semantic segmentation" refers to spatially-dense discrete classification over an area. We group 11 classification evaluations into thematic mapping applications including land use, land cover, crop detection, crop type, and species distribution mapping. These classification datasets vary in their number of classes, complexity of semantics they represent, and their summary periods, i.e., instantaneous observations versus persistence over a reference period (Table 1). Assessing max-trial performance in Figure 1A, we find that AEF achieves the greatest error reductions for evaluations over annual periods: e.g., LCMAP land cover, Descals oil palm, Africa crop mask, and LCMAP land use. Other than Ethiopia crops, all thematic mapping evaluations had > 1.0x reductions in error within the $\sim$90confidence interval. AEF’s consistent performance across these diverse evaluations suggests a degree of generality that was previously not possible even with higher-dimensional learned embeddings.
@@ 6 Estimating biophysical variables
The estimation of biophysical variables goes beyond problems of semantics and perception. Effectively extrapolating sparse measurements of
{strip}
Figure 3: Detailed quantitative and qualitative results from select evaluations. The black dotted-line indicates random chance for classification evaluations. Error bars indicate $1\sigma$accuracy / R2 or $\sim$68.27confidence interval by bootstrapping and k-folds when possible. Most baselines are completely unable to explain evapotranspiration from our testing, achieving a negative $R^{2}$, and so the
properties not easily observed in satellite or other overhead imagery stands to benefit applications from greenhouse gas emissions to the heating/cooling impact of crops. We consider two biophysical variables: emissivity, which is a unitless measurement of surface radiation, and evapotransportation, which characterizes loss of water to the atmosphere from Earth’s land surface. In the max-trial setting, we find that all baselines were able to explain emissivity for some method of transfer with $R^{2}$ > 0.5 except for xy, xyz, and CCDC. AEF had the highest $R^{2}$ (0.72±0.000.72$\pm$ 0.000.72 ± 0.00), followed by MOSAIKS (0.69±0.000.69$\pm$ 0.000.69 ± 0.00). In characterizing evapotranspiration, AEF demonstrates a significant departure from the other baselines tested being the only method with $R^{2}>0.2$, achieving $R^{2}=0.58\pm 0.01$ (Figure 3). We note that the two baselines with explanatory power in this evaluation, composites and MOSAIKS, are simple transformations from raw satellite data, indicating a gap in applicability of both learned and designed featurization approaches prior to AEF.
@@ 7 Change detection
[…]
Figure 4: Effects of scaling. Error bars indicate $1\sigma$ accuracy / $R^{2}$ or $\sim$68.27confidence interval by bootstrapping and k-folds when possible. (A) BA as a function of training examples in AEF for select evaluations compared to other learned featurization approaches. AEF generally outperforms other approaches when trained on the same number of unique observations or fewer. From published documentation, SatCLIP uses 100k observations, Prithvi 4.2M observations, and Clay 70M. (B) The effect of compounding training targets on BA for select evaluations. All BA differences for each additional source group are significant for $\alpha=5\%$, though saturation effects are apparent following the LiDAR or Environmental source group for some evaluations.

### v2

@@ Abstract
Unprecedented volumes of Earth observation data are continually collected around the world, but high-quality labels remain scarce given the effort required to make physical measurements and observations. This has led to considerable investment in bespoke modeling efforts translating sparse labels into maps. Here we introduce AlphaEarth Foundations, an embedding field model yielding a highly general, geospatial representation that assimilates spatial, temporal, and measurement contexts across multiple sources, enabling accurate and efficient production of maps and monitoring systems from local to global scales. The embeddings generated by AlphaEarth Foundations are the only to consistently outperform a suite of other well-known/widely accepted featurization approaches tested on a diverse set of mapping evaluations without re-training. We have released a dataset of global, annual, analysis-ready embedding field layers from 2017 through 2024.
@@ 2 From sparse labels to maps
evaluation shown, all available training data or the "max-trial" was used. Error bars indicate the $1\sigma$ best and worst case ratio or $\sim$90% confidence interval by bootstrapping and k-folds when possible. Evaluations are sorted by increasing mean error ratio. The dashed line represents a ratio of 1 with higher values indicating AEF outperforming the next best solution. (B) AEF reconciles multiple sparse, non-uniformly sampled observation records into a continuous record, regardless of fluctuations in availability. AEF "embedding fields" are a result of static temporal summaries drawn over a conditional "valid period" that need not fully intersect the "support period", where the latter defines the temporal range of the input data. Multiple raster and scalar measurement sources are modeled as sources or targets by AEF. These may be any combination of temporally, geographically, and spatially sparse. In the example shown here, NLCD is not present, and GEDI is available in only a sparse fraction of the spatial context. (C) A view of our global embedding field for the year 2023, note apparent climatic gradients at large scales. (D) AEF produces highly resolved features at 10m2, shown here plotting arbitrary axes in Oaxaca, Mexico. (E) A stack of 64 rasterized AEF layers forms an embedding field, and each individual vector maps to a coordinate on the unit sphere $S^{63}$.
@@ 3 AlphaEarth Foundations
AlphaEarth Foundations (AEF) is the only learned EO featurization approach to outperform a representative sample of featurization methods tested across a broad set of sparse data domains (Figure 1A), reducing error magnitudes by $\sim$23.9% ($\sim$1.4x error magnitude reduction) on average while maintaining a best-in-class 10-meter spatial resolution that requires 16x less information per-representation (64 bytes) compared to the next-most compact learned method. We achieve this leap in performance across challenging mapping applications through a number of innovations; namely, we employ an adaptive decoding scheme that considers time and sensor parameters as continuous variables in an implicit decoder with associated losses, a spatially dense information time-bottleneck, time-conditional summarization, and spatially-precise alignment with geotagged text from Wikipedia articles joined with locations from Global Biodiversity Information Facility (GBIF) species occurrence records. To the best of our knowledge, AEF is the first EO featurization approach to support continuous time (Figure 1B; see supplemental materials S16.2.1). Additionally we introduce a challenging evaluation suite composed of high-quality reference data that attempts to faithfully replicate realistic mapping scenarios. We make our annualized planet-scale feature maps or "embedding field" layers (Figure 1C-D) and evaluation suite available under an open license to encourage further exploration and use.
@@
AEF is designed to accept $N_{i}$ frames for $i\in M_{E}$ input (encoded) data sources with $C_{i}$ channels resampled to the same spatial resolution, and a millisecond epoch timestamp $t_{j},1\leq j\leq{\Sigma}N_{i}$ (Figure 2A). The range of the input timestamps we refer to as the “support period”. For the purposes of learning or at inference time, we support a pair of conditioning timestamps or “valid period” $t_{s},t_{e}$ where $t_{s}<t_{e}$ provides a temporal summary of the Earth’s surface and climatic activity over $t_{s},t_{e})$, even when there is no $t_{j}$ for which $t_{s}\leq t_{j}<t_{e}$ (interpolation), or when $t_{e}\leq t_{j},\forall j\in\{1,2,...,{\Sigma}N_{i}\}$ or $t_{s}>t_{j},\forall j\in\{1,2,...,{\Sigma}N_{i}\}$ (extrapolation). These summaries or “embeddings” are 64 bytes in size, and each embedding contains information that reproduces the temporal trajectory of variables listed in Table S1 over the summary period (Figure 2B) using conditional metadata from each source (see supplemental materials S16.2.1). By explicitly separating the input intervals from those used for the temporal summary, we can apply AEF to time dependent problems requiring a precise date range without fine-tuning. Embeddings are further constrained to distribute uniformly in $S^{63}$ using a so-called “batch uniformity” objective (Figure 2C; see supplemental materials S16.2.4).
@@
Figure 2: AlphaEarth Foundations. (A) Block diagram of the overall network architecture used for video analysis. Preprocessing converts raw observation data via normalization using global statistics, and acquisition timestamps are converted to sinusoidal timecodes. Individual source encoders transform inputs to the same latent space before entering the bulk of the model. Outputs are summarized using conditional timecodes or "summary periods", unique to each decoded source and contrastive learning task. $\mu$ refers to the embedding outputs of the model. The grey shaded region constitutes the model. (B) Model outputs are treated as the mean direction of a von Mises-Fisher distribution, and decoding proceeds by sampling this distribution, and concatenating it with sensor geometry metadata and a timecode indicating the relative position in the valid period to decode. Decoding proceeds for all sources, with losses dependent on the characteristics of each source (see supplemental materials S15). (C) To prevent collapse and improve performance, embeddings are compared to equivalent batch-rotated embeddings using a dot product. The absolute value of this quantity is minimized as a necessary condition for an empirically uniform distribution in $S^{63}$. (D) Block diagram of the model bulk, consisting of simultaneous pathways at different resolutions to maintain efficiency and spatial precision. (E) Contrastive learning between the video teacher and student model, and text encoder. (F) Complete 360° view of 2023 annual embedding field covering Earth’s land surface including minor islands over approximately $\pm$ 82°.
@@
Our video summarization architecture must simultaneously maintain highly localized representations as well as model long distance relationships through time and space in a computationally efficient way; for this we’ve designed an encoder termed Space Time Precision or “STP” that consists of repeated blocks of three simultaneous operators interleaved with spatial pyramid "exchanges" (Figure 2D) inspired by Wang et al. (2020) but more efficiently utilizing learned resampling stages. Given a square input of $L$ pixels a side, each block consists of a $\frac{1}{16}L$ "space" operator following ViT-like spatial self-attention (Dosovitskiy et al., 2020), a $\frac{1}{8}L$ “time” operator utilizing time-axial self-attention, and a $\frac{1}{2}L$ "precision" operator utilizing 3x3 convolutions. Each sequence element in the “time” operator is conditioned on the associated input timestamp ($t_{j}$) after conversion to a sinusoidal timecode. STP blocks terminate with learned Laplacian pyramid rescaling, such that each operator can pass its state to each of the operators in the subsequent block. STP itself terminates with a final learned spatial resampling to the resolution of the "precision" operator. Thus for ${\sum}N_{i}$ inputs, STP produces ${\sum}N_{i}$ output feature maps at $\frac{1}{2}L$ pixels resolution. Initial down-scaling to $\frac{1}{2}L$ is performed in the input projectors marked with an "H" in Figure 2A.
@@ 4 Evaluation in realistic data-scarce scenarios
To establish the performance of AEF relative to other learned and designed representations, we required an evaluation dataset that included archetypal examples of realistic mapping applications. These include thematic mapping, biophysical variable estimation, and change detection at annual and sub-annual cadences. We found most if not all datasets in existing geospatial benchmark suites provide annotations at an object- or image-level rather than pixel-level, rely on labels sampled from existing (machine-generated) datasets, require running the benchmark analysis using provided source imagery, have limited geographic coverage, and/or do not provide sufficient spatial precision or temporal information, e.g., (Lacoste et al., 2023; Bountos et al., 2023), limiting their value in assessing practical use.
To address the need for high-quality labeled datasets that can be used to simulate low-shot (i.e., tens to hundreds of samples) performance in data-scarce regimes requiring precise map outputs, we developed a set of 15 evaluations sourced from 11 openly available datasets. These datasets were selected to represent archetypal classification, regression, and change detection use cases, with all datasets directly linked to real-world products and applications, including land use/land cover mapping and change detection, crop type mapping at different hierarchies, tree genera and plantation classification, and estimating evapotranspiration and emissivity at national to global scales (Table 1). For each evaluation dataset, we selected a balanced number of training samples from each class or equally-spaced partition, with the number of samples determined based on the minimum class size (Table 1, Max Trial Size) and the remainder of the dataset reserved for testing. We then assess performance across a suite of trials designed to test very-low shot training sizes with $n$ samples per class ($n=1$, $10$, max) via transferal methods with minimal parameters: k-nearest neighbors, and linear layers fit to the features (see supplemental materials S18). The “max-trial” scenario is meant to represent more realistic sparse dataset sizes (hundreds as opposed to thousands or millions of points), whereas the ten and one shot trials are meant to evaluate performance given extreme data sparsity.
@@
We used this set of evaluations to compare AEF with a representative set of domain-specific baselines specifically designed for Earth observation applications, including three designed featurization approaches: CCDC (Zhu and Woodcock, 2014; Gorelick et al., 2023), MOSAIKS (Rolf et al., 2021), and composites (Qiu et al., 2023), and three learned featurization approaches: SatCLIP (Klemmer et al., 2025), Prithvi (Jakubik et al., 2023b), and Clay (Clay, 2024). We also include three controls: spatial coordinates (XY), coordinates and elevation (XYZ), and a ViT (Vision Transformer) pre-trained on ImageNet (Dosovitskiy et al., 2020; Deng et al., 2009). Where applicable, baselines were provided with identical inputs to AEF and baseline hyperparameters were tuned to maximize performance on our evaluation suite (see supplemental materials S19).
Our evaluations showed AEF consistently outperforms both designed and learned featurization methods in all trial settings. AEF reduced error magnitudes overall by $\sim$23.9% on average when compared to the next-best approach and method of transfer in the max-trial setting (Figure 1A). For ten-shot trials, AEF reduced error magnitudes by $\sim$10.4% on average compared to the next-best approach, and for one-shot trials AEF reduced error magnitudes by $\sim$4.18%. We show quantitative and qualitative results for select evaluations in Figure 3 (and see supplemental materials S20 for full quantitative results). The next-best approach varies across evaluation dataset and method, indicating both non-uniform progress and that AEF unlocks progress in historically challenging mapping scenarios. We discuss these results and the effect of scaling training data in further detail below.
@@ 5 Thematic mapping
Thematic mapping or "semantic segmentation" refers to spatially-dense discrete classification over an area. We group 11 classification evaluations into thematic mapping applications including land use, land cover, crop detection, crop type, and species distribution mapping. These classification datasets vary in their number of classes, complexity of semantics they represent, and their summary periods, i.e., instantaneous observations versus persistence over a reference period (Table 1). Assessing max-trial performance in Figure 1A, we find that AEF achieves the greatest error reductions for evaluations over annual periods: e.g., LCMAP land cover, Descals oil palm, Africa crop mask, and LCMAP land use. Other
{strip}
Figure 3: Detailed quantitative and qualitative results from select evaluations. The black dotted-line indicates random chance for classification evaluations. Error bars indicate $1\sigma$ accuracy / R2 or $\sim$68.27% confidence interval by bootstrapping and k-folds when possible. Most baselines are completely unable to explain evapotranspiration from our testing, achieving a negative $R^{2}$, and so the
than Ethiopia crops, all thematic mapping evaluations had > 1.0x reductions in error within the $\sim$90% confidence interval. AEF’s consistent performance across these diverse evaluations suggests a degree of generality that was previously not possible even with higher-dimensional learned embeddings.
@@ 6 Estimating biophysical variables
The estimation of biophysical variables goes beyond problems of semantics and perception. Effectively extrapolating sparse measurements of properties not easily observed in satellite or other overhead imagery stands to benefit applications from greenhouse gas emissions to the heating/cooling impact of crops. We consider two biophysical variables: emissivity, which is a unitless measurement of surface radiation, and evapotranspiration, which characterizes loss of water to the atmosphere from Earth’s land surface. In the max-trial setting, we find that all baselines were able to explain emissivity for some method of transfer with $R^{2}$ > 0.5 except for xy, xyz, and CCDC. AEF had the highest $R^{2}$ ($0.72\pm 0.00$), followed by MOSAIKS ($0.69\pm 0.00$). In characterizing evapotranspiration, AEF demonstrates a significant departure from the other baselines tested being the only method with $R^{2}>0.2$, achieving $R^{2}=0.58\pm 0.01$ (Figure 3). We note that the two baselines with explanatory power in this evaluation, composites and MOSAIKS, are simple transformations from raw satellite data, indicating a gap in applicability of both learned and designed featurization approaches prior to AEF.
@@ 7 Change detection
In the max-trial setting with direct supervision of change, we find that AEF’s performance exceeds performance of other models and datasets achieving $78.4\%\pm 1.11$ balanced accuracy (BA) (linear) and $79.3\%\pm 1.67$ BA (kNN, k=3) on the land cover and land use evaluations respectively. The next-best baseline achieved $72.0\%\pm 1.28$ BA (MOSAIKS, kNN, k=3) and $71.5\%\pm 2.33$ BA (composite, kNN, k=3) respectively. In the max-trial setting with unsupervised thresholding, AEF exceeds performance of all other baselines when detecting land cover change (Figure 3), but not land use change, achieving $71.3\%\pm 1.14$ BA and $71.4\%\pm 2.08$ BA compared to $67.0\%\pm 1.28$ BA (ViT) and $72.9\%\pm 1.97$ BA (ViT) respectively, suggesting the value of supervision for this use case.
Figure 4: Effects of scaling. Error bars indicate $1\sigma$ accuracy / $R^{2}$ or $\sim$68.27% confidence interval by bootstrapping and k-folds when possible. (A) BA as a function of training examples in AEF for select evaluations compared to other learned featurization approaches. AEF generally outperforms other approaches when trained on the same number of unique observations or fewer. From published documentation, SatCLIP uses 100k observations, Prithvi 4.2M observations, and Clay 70M. (B) The effect of compounding training targets on BA for select evaluations. All BA differences for each additional source group are significant for $\alpha=5\%$, though saturation effects are apparent following the LiDAR or Environmental source group for some evaluations.